OpenStudy (anonymous):

So apparently x^2 + 2^2 = 3^2 Is: +-sqrt(5) How do you get to this? I get +-sqrt(23)

OpenStudy (anonymous):

\[x^2 + 2^2 = 3^2\] \[x^2=3^2 - 2^2\] \[x^2 = 9- 4\] \[x^2 = 5\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (phi):

how do you get sqrt(23) ? if you immediately change 2^2 = 2*2 to 4 and 3*3 to 9 you have x^2 +4= 9

OpenStudy (anonymous):

I'm an idiot....... :)

OpenStudy (anonymous):

thanks!